OpenStudy (anonymous):

quadratic equations: solve: y(y+1) = 2?

jhonyy9 (jhonyy9):

y2 +y-2=0 (y+2)(y-1)=0

OpenStudy (turingtest):

multiply out the y into the parentheses, then rearrange the equation to be of the form\[ax^2+bx+c=0\]then factor, use the quadratic formula, or complete the square